# Griesmer and Optimal Linear Codes from the Affine Solomon-Stiffler Construction

### v1

@@ 8 Conclusion
Two-weight $q$-ary Griesmer and optimal linear codes have been constructed in two published papers [26, 18] in IEEE Transactions on Information Theory. Their weight distributions were determined. In this paper, we propose a general construction of Griesmer and optimal linear codes as an affine version of Solomon-Stiffler codes. Modified affine Solomon-Stiffler codes include these projective Solomon-Stiffler codes as special cases. Codes in [26, 18] have the same parameters and weight distributions as very special case codes in our construction. Our Griesmer and optimal codes are less restrictive than Griesmer and optimal trace codes constructed in [15]. Optimal binary codes constructed in a recent paper [24] and another paper [25] and codimension one subcodes of special binary Solomon-Stiffler codes have the same parameters. We reconstructed many optimal codes in [9] from our modified affine Solomon-Stiffler construction. Projective linear almost optimal codes similar to codes in [28] were also constructed.

### v2

@@ 9 Conclusion
In this paper, we propose a general construction of Griesmer and optimal linear codes as an affine version of Solomon-Stiffler codes. Modified affine Solomon-Stiffler codes include these projective Solomon-Stiffler codes as special cases. Griesmer or optimal codes constructed in [27, 19] have the same parameters and weight distributions as very special case codes in our construction. Our Griesmer and optimal codes are less restrictive than Griesmer and optimal trace codes constructed in [16]. Optimal binary codes constructed in a recent paper [25] and another paper [26] and codimension one subcodes of special binary Solomon-Stiffler codes have the same parameters. Distance-optimal codes as punctured code of Griesmer codes were also discussed. We reconstructed many optimal codes in [9] from our modified affine Solomon-Stiffler construction. Projective linear almost optimal codes similar to codes in [30] were also constructed.